OpenStudy (mary.rojas):

Given a parabola with zeros at x=0 and x=4 , which of the following MUST be true: I. The parabola opens upward II. The parabola has a line of symmetry at x=2 III. The parabola intercepts the y-axis at y=0 a. I only b. III only c. I and III only d. II and III only e. I, II, and III

OpenStudy (anonymous):

The parabola could open either up or down: \[y = (x-0)(x-4)=x^2-4x \textrm{ or } y = -(x-0)(x-4) = -x^2 + 4x\] So it can't be \(\textrm{A, C,}\) or \(\textrm{E}\). Also, the axis of symmetry is always the midpoint of the two roots (or zeros, or x-intercepts). So \[x_{sym} = \dfrac{0+4}{2} = 2\] means that \(\textrm{B}\) is false, leaving only \(\textrm{D}\). And if you're curious about \(\textrm{III}\): Zeros are points on a graph where \(y=0\). At \(x = 0\) (The \(y\)-axis), \(y\) also equals 0, meaning that the graph intercepts the y-axis at \(y=0\), or \((0,0)\).

OpenStudy (mary.rojas):

thank you!!!!! you are a life saver!!!